Hope Industries just paid a dividend of $2.00 per share (i.e., D0 = $2.00). Analysts expect the company's dividend to grow 40 percent this year, and 20 percent in second year. After two years the dividend is expected to grow at a constant rate of 6 percent. The risk free rate is 4% and expected market risk premium is 6% and the firm is twice as risky as market. First calculate the current stock price using Excel. If the target price of the company's stock is $50.00, what should be the expected stable constant growth rate after two years? You must use “Goal Seek” command in Excel
Answer #1

Calculation of current stock price is as follows

Expected return Ke = risk free rate + Market risk premium * Beta

Given firm is twice risky as market, therefore beta is 0.06*2 = 1.2

Expected return = 0.04 + 0.06 * 1.2 = 0.112 = 11.20%

Current stock price P0 = D0 (1+g)/ Ke - g = 2 (1+0.06)/0.112-0.06 = $ 40.77

Calculation of stable growth rate is as follows:

50 = 2.12(1+g)/(0.112-g) = g = 0.0726 = 7.26%